# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ Abstract.
Software defect datasets, which are collections of software bugs and their associated information, are essential resources for researchers and practitioners in software engineering and beyond. Such datasets facilitate empirical research and enable standardized benchmarking for a wide range of techniques, including fault detection, fault localization, test generation, test prioritization, automated program repair, and emerging areas like agentic AI-based software development. Over the years, numerous software defect datasets with diverse characteristics have been developed, providing rich resources for the community, yet making it increasingly difficult to navigate the landscape. To address this challenge, this article provides a comprehensive survey of 132 software defect datasets. The survey discusses the scope of existing datasets, e.g., regarding the application domain of the buggy software, the types of defects, and the programming languages used. We also examine the construction of these datasets, including the data sources and construction methods employed. Furthermore, we assess the availability and usability of the datasets, validating their availability and examining how defects are presented. To better understand the practical uses of these datasets, we analyze the publications that cite them, revealing that the primary use cases are evaluations of new techniques and empirical research. Based on our comprehensive review of the existing datasets, this paper suggests potential opportunities for future research, including addressing underrepresented kinds of defects, enhancing availability and usability through better dataset organization, and developing more efficient strategies for dataset construction and maintenance.
@@ 1. Introduction
Software defect datasets are curated collections of real-world software bugs. These datasets serve as standardized benchmarks, fostering consistent and reproducible evaluations of novel techniques across diverse studies. Additionally, the realism, diversity, scale, and accessibility of these datasets support empirical analyses of defect characteristics and enable the training and fine-tuning of machine learning models for software quality assurance. For instance, the widely adopted Defects4J dataset (Just et al., 2014) has been used for a wide range of technical evaluations (e.g., (Martinez et al., 2017; Zhong et al., 2025; Ouyang et al., 2024)), empirical studies (e.g., (Sobreira et al., 2018; Durieux et al., 2019)), and to train machine learning models (e.g., (Zhang et al., 2024c)). As another example, the recently proposed SWE-Bench (Jimenez et al., 2024) dataset is widely used to evaluate the effectiveness of large language models (LLMs) in automating software engineering (Zhang et al., 2024d; Yang et al., 2024) and serves as the basis for other benchmarks (Aleithan et al., 2024; Mündler et al., 2024), demonstrating the practical value of software defect datasets.
With the continuous growth in the complexity and scale of modern software systems, there is an increasing demand for comprehensive and high-quality defect datasets. This demand is further amplified by the rising popularity of artificial intelligence (AI)-based software engineering approaches. In response to this escalating demand, industry challenges, competitions, and workshops have been organized to encourage the development of software defect datasets of high quality. As a result, these datasets have experienced a significant increase in both quantity and diversity in recent years, becoming essential resources for researchers and practitioners in multiple fields, including software engineering, programming languages, computer security, computer systems, and artificial intelligence.
@@
To address these challenges, this paper provides a comprehensive overview of existing software defect datasets. Specifically, we follow a systematic methodology (detailed in Section 2) for collecting and filtering relevant literature, resulting in the identification of 132 software defect datasets. Our survey focuses on the five key aspects of software defect datasets, as illustrated in Figure 1:
(1) Scope of the datasets
@@
For each dataset surveyed, we examine the application domain of the software systems from which the defects are collected (e.g., machine learning systems or mobile applications), the types of defects (e.g., functional, security, or performance), the programming languages (e.g., Java, C/C++, or Python), and other characteristics that define the focus of the dataset.
(2) Construction of the datasets
@@
For each dataset, we investigate how the dataset is constructed. This includes the sources of defects (e.g., issue reports, version control systems, or continuous integration & delivery) and the methodologies used to obtain the defects (e.g., manual inspection, keyword matching, or machine learning techniques).
(3) Availability and usability of the datasets
@@
For each dataset, we evaluate its availability and usability. Availability refers to whether the dataset is publicly available, and if so, where it is hosted. Usability refers to how the defects are presented in the dataset, such as providing only textual descriptions, code snippets, or executable artifacts to reproduce the defects.
(4) Uses of the datasets
Focusing on the five most cited datasets in this survey, we conduct a retrospective citation analysis detailing how these datasets have been utilized by the research community. This analysis provides insights into the usage trends over time, highlighting the impact of these datasets on the research community.
(5) Opportunities for future research
@@
Based on our analysis of existing datasets, we identify potential opportunities for future research, including automating the mining of high-quality datasets, e.g., by leveraging CI pipelines and LLMs. Additionally, we identify the need for specialized datasets targeting specific domains, defect types, and new languages, as well as techniques ensuring dataset reproducibility and continuous updates.
@@
Our work contributes the first comprehensive survey on software defect datasets, providing a detailed discussion of the scope, construction, availability, and real-world usage of datasets. Orthogonal to existing surveys on software defects, which focus on categorizing (Nagwani and Verma, 2014; Zhang et al., 2015), detecting/localizing (Wong et al., 2016; Lin et al., 2020; Paltenghi and Pradel, 2024; Chen et al., 2021a; Andreasen et al., 2017), and repairing software defects (Monperrus, 2018; Huang et al., 2023; Zhang et al., 2024b) without collecting them, this survey paper focuses on the datasets that collect software defects that are prevalent in software codebases. Also, previous literature reviews on 73 benchmarks (Hirsch and Hofer, 2022) or 22 bug datasets (Holek et al., 2023) provide a high-level overview of the benchmarking datasets on the aspects like size, type, and programming languages. In contrast, this article offers significantly broader coverage of 132 software defect datasets, with a more in-depth analysis of their scope, construction, availability, usability, and practical uses.
@@ 2. Survey Methodology
This survey is conducted on 132 software defect datasets. To identify these datasets, we follow a systematic process that collects papers from DBLP [^2] https://dblp.org/ and selects papers through manual review. For paper collection, we search through DBLP with two complementary search strategies, summarized in Table 1, to ensure comprehensive coverage of research on software defect datasets.
Table 1. Searching strategies to collect papers from DBLP.
Strategy Search Content #Papers
@@
Key Phrase bug benchmark, bug dataset, defect benchmark, defect dataset, fail pass, fault benchmark, fault dataset, mining software, software benchmark, software bug, software dataset, software defect, software fault, software vulnerabilit(y/ies), vulnerabilit(y/ies) benchmark, vulnerabilit(y/ies) dataset, empirical bug, empirical defect, empirical error, empirical failure, empirical fault, empirical issue, study bug, study defect, study error, study failure, study fault, study issue 11,280
@@
Keyword benchmark, bug, dataset, defect, fault, issue, vulnerabilit(y/ies) 11,872
@@
Venue ASE, FSE, ICSE, ISSRE, ISSTA, MSR, OOPSLA, SANER, TOSEM, TSE, AAAI, ACL, EMNLP, ICLR, ICML, NeurIPS
(1) Key Phrase
We use specific key phrases to identify papers with titles explicitly related to software defect datasets or empirical studies on software defects. These key phrases help us target work that directly focuses on collecting or analyzing software defect data. Papers with titles containing all words in a key phrase are included, resulting in 11,280 papers.
(2) Keyword ×Venue
To complement the key phrase-based search results, we also use a set of individual keywords. As searching with individual keywords yields a very large number of papers, we restrict this search to high-impact venues in software engineering and machine learning. These selected software engineering venues are known for publishing a substantial amount of research on software defects, and machine learning venues are included to capture studies at the intersection of machine learning and software defect research. From the selected venues, we include papers containing at least one of the keywords, resulting in 11,872 papers.
We set the cutoff date for paper selection as January 1, 2025. Only papers that have been peer-reviewed and published before the cutoff date are included, thus excluding preprints. After de-duplicating papers found by both search strategies, we identified a total of 15,777 papers potentially related to software defect datasets, which we proceeded to check manually by reviewing their titles, abstracts, and full texts.
One author of this survey conducted an initial review of the titles of all 15,777 papers and selected a subset of more than 300 candidates that appeared likely to introduce software defect datasets. Then, two authors carefully reviewed and analyzed these candidates, jointly deciding to discard those that did not meet our inclusion criteria. We select a paper if its main contribution is one or more software defect datasets, or if it includes software defects as byproducts. Papers are excluded if they only used existing datasets without introducing new ones, presented datasets for purposes unrelated to software defects (e.g., code generation or performance benchmarking), or focused solely on techniques for mining, injecting, or labeling defects. After reaching agreement on all included datasets, we ended up with 132 software defect datasets from 131 papers (one paper (Le Goues et al., 2015) presents two distinct datasets). The manual review process spanned several months.
@@
3. Scope of Software Defect Datasets
@@ 3.1. Application Domains of Defective Software
Machine Learning Applications (Morovati et al., 2023; Liang et al., 2022; Kim et al., 2021b; Zhang et al., 2018; Chen et al., 2021b; Wang et al., 2022b) 6
Frameworks (Guan et al., 2023; Liu et al., 2022; Ho et al., 2023; Zhao et al., 2023b) 4
Compilers (Du et al., 2021; Shen et al., 2021) 2
Compilers & Runtime Engines (Zhang et al., 2024a; Yu et al., 2024; Yu and Wang, 2024; Paul et al., 2021; Du et al., 2021; Zaman et al., 2012; Xu et al., 2023; Wang et al., 2023a, c; Shen et al., 2021) 10
Mobile Applications (Tebib et al., 2024; Das et al., 2024; Senanayake et al., 2023; Xiong et al., 2023; Wendland et al., 2021; Riganelli et al., 2019; Mitra and Ranganath, 2017; Liu et al., 2019; Chen et al., 2021b; Tan et al., 2018) 10
Blockchain Smart Contracts (Cai et al., 2024; Chen et al., 2022; Zhang et al., 2020; Wang et al., 2023b) 4
Clients (Kim et al., 2022) 1
Autonomous Systems (Timperley et al., 2024; Garcia et al., 2020; Wang et al., 2021a) 3
Quantum Computing Applications (Zhao et al., 2023a) 1
Platforms (Paltenghi and Pradel, 2022; Zhao et al., 2023b) 2
Scientific Computing (Murphy et al., 2020; Azad et al., 2023; Franco et al., 2017) 3
Cloud Computing (Xu et al., 2024; Gunawi et al., 2014) 2
Databases (Liu et al., 2024b; Cui et al., 2024) 2
Computational Notebooks (de Santana et al., 2024) 1
Internet of Things (Makhshari and Mesbah, 2021) 1
@@ 3.1.1. Defects Related to Machine Learning
The reliability and robustness of machine learning applications and their underlying software platforms are crucial for their successful deployment. Machine learning-related software defects can cause severe consequences, such as incorrect predictions, biased decisions, and performance degradation (Zhang et al., 2018). We identify three subdomains of machine learning-related software, each with several defect datasets: machine learning applications, machine learning frameworks, and machine learning compilers.
@@
Defects in machine learning applications often arise during model development, training, and deployment, leading to critical issues like incorrect predictions or silent errors. Datasets such as defect4ML (Morovati et al., 2023), gDefects4DL (Liang et al., 2022), and Denchmark (Kim et al., 2021b) focus on common machine learning defects, including API misuses, performance problems, and numerical errors. There are also machine learning defect datasets provided as by-products of empirical studies, such as studies on bugs from programs relying on TensorFlow (Zhang et al., 2018), on numerical defects in deep learning programs (Wang et al., 2022b), and on deployment-related bugs in machine learning-based mobile applications (Chen et al., 2021b).
@@
Machine learning frameworks, such as PyTorch and TensorFlow, are crucial for building and optimizing models, but themselves may suffer from defects. Software defect datasets for machine learning frameworks typically result from empirical research. For example, MOB-Dataset (Guan et al., 2023) collects real-world machine learning optimization defects from TensorFlow and PyTorch, and comes with an empirical study of model optimization. Likewise, empirical studies of machine learning frameworks have produced defect datasets of common bugs in PyTorch (Ho et al., 2023) and quantum machine learning frameworks (Zhao et al., 2023b). Other work presents a dataset and a defect taxonomy of aging-related bugs in machine learning libraries (Liu et al., 2022). These studies provide critical insights into the defects that can impact the functionality, efficiency, and stability of machine learning frameworks.
@@
In addition to applications and frameworks, machine learning compilers play a vital role in optimizing models, but they also introduce unique bugs. Based on datasets of defects in machine learning compilers, Du et al. (2021) and Shen et al. (2021) both examine frequent issues in these compilers, such as incorrect optimizations, numerical inaccuracies, and incompatibilities with hardware accelerators.

### v2

@@ Abstract.
Software defect datasets, which are collections of software bugs, are essential resources to facilitate empirical research and enable standardized benchmarking for a wide range of software engineering techniques, including emerging areas like agentic AI-based software development. Over the years, numerous software defect datasets have been developed, providing rich resources for the community, yet making it increasingly difficult to navigate the landscape. This article provides a comprehensive survey of 151 software defect datasets, covering their scope, construction, availability, usability, and practical uses. We also suggest potential opportunities for future research based on our findings, such as addressing underrepresented kinds of defects.
@@ 1. Introduction
Software defect datasets are curated collections of real-world software bugs. These datasets serve as standardized benchmarks, fostering consistent and reproducible evaluations of novel techniques across diverse studies. Additionally, the realism, diversity, scale, and accessibility of these datasets support empirical analyses of defect characteristics and enable the training and fine-tuning of machine learning models for software quality assurance. For instance, the widely adopted Defects4J dataset (DBLP:conf/issta/JustJE14) has been used for a wide range of technical evaluations (e.g., (Martinez et al., 2017; Zhong et al., 2025; Ouyang et al., 2024)), empirical studies (e.g., (Sobreira et al., 2018; Durieux et al., 2019)), and to train machine learning models (e.g., (Zhang et al., 2024b)). As another example, the recently proposed SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24) dataset is widely used to evaluate the effectiveness of large language models (LLMs) in automating software engineering (Zhang et al., 2024c; Yang et al., 2024) and serves as the basis for other benchmarks (Aleithan et al., 2024; Mündler et al., 2024), demonstrating the practical value of software defect datasets.
With the continuous growth in the complexity and scale of modern software systems, there is an increasing demand for comprehensive and high-quality defect datasets. This demand is further amplified by the rising popularity of artificial intelligence (AI)-based software engineering approaches (Hou et al., 2024; Yang et al., 2022). In response to this escalating demand, industry challenges, competitions, and workshops have been organized to encourage the development of software defect datasets of high quality. As a result, these datasets have experienced a significant increase in both quantity and diversity in recent years, becoming essential resources for researchers and practitioners in multiple fields, including software engineering, programming languages, computer security, computer systems, and artificial intelligence.
@@
To address these challenges, this paper provides a comprehensive overview of existing software defect datasets. Specifically, we follow a systematic methodology (detailed in Section 2) for collecting and filtering relevant literature, resulting in the identification of 151 software defect datasets. Our survey focuses on the following four research questions (RQs), as illustrated in Figure 1:
@@
  • RQ1: What is the scope of existing software defect datasets? (Section 3) For each dataset surveyed, we examine the application domain of the software systems from which the defects are collected (e.g., machine learning systems or mobile applications), the types of defects (e.g., functional, security, or performance), the programming languages (e.g., Java, C/C++, or Python), and other characteristics that define the focus of the dataset. We also explore potential interrelationships among these dimensions to provide a holistic understanding of the dataset landscape.
@@
  • RQ2: How are existing software defect datasets constructed? (Section 4) For each dataset, we investigate how the dataset is constructed. This includes the sources of defects (e.g., issue reports, version control systems, or continuous integration & delivery) and the methodologies used to obtain the defects (e.g., manual inspection, keyword matching, or machine learning techniques).
@@
  • RQ3: To what extent are existing software defect datasets available and usable in practice? (Section 5) For each dataset, we evaluate its availability and usability. Availability refers to whether the dataset is publicly available, and if so, where it is hosted. Usability refers to how the defects are presented in the dataset, such as providing only textual descriptions, code snippets, or executable artifacts to reproduce the defects.
  • RQ4: How are software defect datasets actually used by research community? (Section 6) While prior literature reviews (Hirsch and Hofer, 2022; Holek et al., 2023) have provided high-level summaries of the intended uses of software defect datasets from the perspective of their creators, this research question seeks to offer a more comprehensive and detailed understanding of how these datasets are actually used. By analyzing citations of the most influential datasets, we aim to uncover usage patterns and trends over time, thereby revealing their impact on the research community.
@@
Finally, based on our analysis of existing datasets, Section 7 identifies potential opportunities for future research, including automating the mining of high-quality datasets, e.g., by leveraging CI pipelines and LLMs. Additionally, we identify the need for specialized datasets targeting specific domains, defect types, and new languages, as well as techniques ensuring dataset reproducibility and continuous updates.
Our work contributes the first comprehensive survey on software defect datasets, providing a detailed discussion of their scope, construction, availability, and real-world usage. Orthogonal to existing surveys on software defects, which focus on categorizing (Nagwani and Verma, 2014; Zhang et al., 2015), detecting/localizing (Wong et al., 2016; Lin et al., 2020; Paltenghi and Pradel, 2024; Chen et al., 2021; Andreasen et al., 2017), and repairing software defects (Monperrus, 2018; Huang et al., 2023; Zhang et al., 2024a), this survey focuses on datasets that collect software defects prevalent in software codebases. Previous literature reviews on 73 benchmarks (Hirsch and Hofer, 2022) and 14 bug datasets (Holek et al., 2023) provide high-level overviews of existing benchmarking datasets. However, as summarized in Table 1, these studies do not cover several important aspects of software defect datasets. In contrast, this article provides (1) significantly broader coverage of 151 software defect datasets with a more recent cutoff date, (2) a systematic and in-depth analysis of their scope, construction, availability, and usability, (3) insights into the relationships between different dataset characteristics, and (4) an examination of longitudinal trends in the actual usage of the most influential datasets.
@@
Table 1. Comparison between surveys of software defect datasets.
Hirsch and Hofer (2022) Holek et al. (2023) This Survey
Number of Datasets 73 14 151
Cutoff Date For Dataset Selection November 2021 April 2022 January 2025
Scope of Datasets
- Application Domains
- Types of Defects
- Programming Languages
Dataset Construction
- Source of Defects
- Mining Methodology
Availability and Usability
- Availability and Hosting Platforms
- Presentation
Uses of Datasets
@@ 2. Survey Methodology
Table 2. Searching strategies to collect papers
Strategy Search Content
@@
Key Phrase bug benchmark, bug dataset, defect benchmark, defect dataset, patch benchmark, patch dataset, fail pass, fault benchmark, fault dataset, mining software, software benchmark, software bug, software dataset, software defect, software fault, software patch, software vulnerabilit(y/ies), vulnerabilit(y/ies) benchmark, vulnerabilit(y/ies) dataset, empirical bug, empirical defect, empirical error, empirical failure, empirical fault, empirical issue, empirical patch, study bug, study defect, study error, study failure, study fault, study issue, study patch
@@
Keyword benchmark, bug, dataset, defect, fault, issue, patch, vulnerabilit(y/ies)
@@
Venue ASE, ESE, FSE, ICSE, ISSRE, ISSTA, MSR, OOPSLA, SANER, TOSEM, TSE, ESEC, AAAI, ACL, EMNLP, ICLR, ICML, NeurIPS, USENIX Security, IEEE S&P, CCS, NDSS
Figure 2. Collection and selection process for software defect dataset papers.
This section describes the methodology used in this survey to identify, select, and analyze existing software defect datasets, guided by research questions described in the introduction. Our survey covers 151 software defect datasets. To identify these datasets, we follow a systematic process that collects papers from multiple bibliographic databases and selects papers through a rigorous manual review. Figure 2 illustrates the complete paper collection and selection process. We conduct a comprehensive literature search across four major bibliographic databases: ACM Digital Library, DBLP, IEEE Xplore, and ScienceDirect. For each source, we execute two complementary search strategies, summarized in Table 2. We include papers that have been peer-reviewed and published before January 1, 2025, excluding preprints.
(1) Key Phrase-Based Search
We use specific key phrases to identify papers explicitly related to software defect datasets or empirical studies on software defects. We search for papers containing all words from a key phrase in their titles, abstracts, and keywords if available. This strategy yields 823 papers from the ACM Digital Library, 10,866 papers from DBLP, 9,107 papers from IEEE Xplore, and 276 papers from ScienceDirect.
(2) Keyword ×Venue Search
To complement the key phrase-based search, we use individual keywords combined with venue restrictions. As individual keywords yield many papers, we restrict this search to high-impact venues in software engineering, machine learning, and software security. We search for papers containing at least one keyword in their title, abstract, and keywords if available. This strategy yields 431 papers from the ACM Digital Library, 6,402 papers from DBLP, and 1,299 papers from IEEE Xplore. Due to API query constraints, this strategy was not applicable to ScienceDirect.
After merging results from both search strategies across all bibliographic databases and performing cross-source deduplication, we obtain 23,057 unique papers potentially related to software defect datasets. Two authors independently review the titles and abstracts of these papers to identify those potentially introducing a software defect dataset. The authors then compare their selections and reach consensus through discussion, excluding 22,728 irrelevant papers and identifying 329 papers for full-text assessment. For the full-text assessment phase, the same two authors independently review each of the 329 papers to determine their eligibility for inclusion. A paper is included if its main contribution is one or more software defect datasets, or if it includes software defects as byproducts. Papers are excluded if they only use existing datasets without introducing a new dataset, present datasets for purposes unrelated to software defects (e.g., code generation or performance benchmarking), or focus solely on techniques for mining, injecting, or labeling defects. The inter-rater agreement between the two authors for the full-text assessment is almost perfect, with a Cohen’s kappa (Cohen, 1960) of 0.871 (95% CI: 0.818–0.924). The authors agreed on 308 out of 329 papers (93.6%), and disagreements on the remaining 21 papers were resolved through discussion and, when necessary, consultation with a third author. This rigorous assessment ultimately leads to 150 papers, from which we identified 151 software defect datasets (one paper (DBLP:journals/tse/GouesHSBDFW15) presents two datasets).
Finally, the same two authors conduct an in-depth manual review of all included datasets to extract and annotate their attributes (e.g., application domain, defect type, and programming language for RQ1; sources of defects and construction methodologies for RQ2; and availability and usability for RQ3). Each dataset is independently cross-checked by both authors to ensure consistency, with ambiguities resolved through discussion. The annotation process is performed manually over several months.
@@
3. Scope of Software Defect Datasets (RQ1)
@@ 3.1. Application Domains of Defective Software
Machine Learning Applications (DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL21; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/kbse/WangWCCY22; DBLP:conf/icse/HumbatovaJBR0T20) 7
Frameworks (DBLP:conf/icse/GuanXLLB23; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:journals/ese/TambonNAKA24) 5
Compilers (DBLP:conf/issre/Du0MZ21; DBLP:conf/sigsoft/ShenM0TCC21) 2
Compilers & Runtime Engines (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/qrs/YuW24; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/msr/ZamanAH12; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/RomanoLK021) 11
Mobile Applications (DBLP:conf/pst/TebibAAG24; DBLP:conf/kbse/DasAM24; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/promise/MitraR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icse/TanDGR18) 10
Blockchain Smart Contracts (DBLP:journals/jss/CaiLZZS24; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icsm/ZhangXL20; DBLP:journals/jss/WangCHZBZ23) 4
Clients (DBLP:conf/msr/KimKL22) 1
Cloud Computing (DBLP:conf/issta/XuG024; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:journals/pacmpl/DrososSAM024) 4
Autonomous Systems (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/sigsoft/WangLX0S21) 3
Quantum Computing Applications (DBLP:journals/jss/ZhaoMLZ23) 1
Platforms (DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/qsw/ZhaoWLLZ23) 2
Scientific Computing (DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/FrancoGR17) 3
Databases (DBLP:journals/tosem/LiuMC24; DBLP:conf/icse/CuiD0WSZWYXH0024) 2
Computational Notebooks (DBLP:journals/tosem/SantanaNAA24) 1
Internet of Things (DBLP:conf/icse/Makhshari021) 1
For clarity, only datasets that explicitly specify the corresponding attribute (i.e., application domain) are listed. The complete list of 151 datasets is available at https://defect-datasets.github.io/ and in the supplementary materials (Zhu et al., 2025). The same convention applies to all tables in Sections 3 and 4.
The application domain of defective software refers to the specific field or area in which the software operates, such as machine learning or cloud computing. Of the 151 software defect datasets considered here, the majority (over $60\%$) focus on general software development and are not tied to any specific application domain. Among the remaining datasets that explicitly target specific domains, we identify eleven distinct application domains (Table 3), including machine learning, compilers and runtime engines, mobile applications, and blockchain. [^2] In this and other classifications of datasets, a single dataset may appear in multiple categories. Some domains attract more attention from dataset creators than others. For example, there are 14 datasets related to machine learning defects, which we further categorize into machine learning applications, machine learning frameworks, and machine learning compilers. The rest of this section discusses the application domains focused on by the software defect datasets, highlighting the uniqueness of defects in each domain.
@@ 3.1.1. Defects Related to Machine Learning
The reliability and robustness of machine learning applications and their underlying software platforms are crucial for their successful deployment. Machine learning-related software defects can cause severe consequences, such as incorrect predictions, biased decisions, and performance degradation (DBLP:conf/issta/ZhangCCXZ18). We identify three subdomains of machine learning-related software, each with several defect datasets: machine learning applications, machine learning frameworks, and machine learning compilers.
@@
Defects in machine learning applications often arise during model development, training, and deployment, leading to critical issues like incorrect predictions or silent errors. Datasets such as defect4ML (DBLP:journals/ese/MorovatiNKJ23), gDefects4DL (DBLP:conf/icse/LiangLSSFD22), and Denchmark (DBLP:conf/msr/KimKL21) focus on common machine learning defects, including API misuses, performance problems, and numerical errors. There are also machine learning defect datasets provided as by-products of empirical studies, such as studies on bugs from programs relying on TensorFlow (DBLP:conf/issta/ZhangCCXZ18), on numerical defects in deep learning programs (DBLP:conf/kbse/WangWCCY22), and on deployment-related bugs in machine learning-based mobile applications (DBLP:conf/icse/ChenYLCLWL21).
@@
Machine learning frameworks, such as PyTorch and TensorFlow, are crucial for building and optimizing models, but themselves may suffer from defects. Software defect datasets for machine learning frameworks typically result from empirical research. For example, MOB-Dataset (DBLP:conf/icse/GuanXLLB23) collects real-world machine learning optimization defects from TensorFlow and PyTorch, and comes with an empirical study of model optimization. Likewise, empirical studies of machine learning frameworks have produced defect datasets of common bugs in PyTorch (DBLP:conf/icsm/HoMI0SKNR23) and quantum machine learning frameworks (DBLP:conf/qsw/ZhaoWLLZ23). Other work presents a dataset and a defect taxonomy of aging-related bugs in machine learning libraries (DBLP:conf/issre/LiuZDHDM022). These studies provide critical insights into the defects that can impact the functionality, efficiency, and stability of machine learning frameworks.
@@
In addition to applications and frameworks, machine learning compilers play a vital role in optimizing models, but they also introduce unique bugs. Based on datasets of defects in machine learning compilers, DBLP:conf/issre/Du0MZ21 and DBLP:conf/sigsoft/ShenM0TCC21 both examine frequent issues in these compilers, such as incorrect optimizations, numerical inaccuracies, and incompatibilities with hardware accelerators.